A student carrying an experiment discovered that it took 2 seconds for a sound wave traveling through a telephone line to cover a distance d meters and 20 seconds for the same sound traveling through air to cover a similar distance. Determine the ratio of speed of sound in air to that in the wire


1
Expert's answer
2021-04-19T07:24:39-0400

By definition, the speed of sound in the wire:


"v_w = \dfrac{d}{2s}"

The speed of sound in air:


"v_a = \dfrac{d}{20s}"

Their ratio is:


"\dfrac{v_a}{v_w} = \dfrac{d\cdot 2s}{20s\cdot d} = \dfrac{2}{20} = \dfrac{1}{10} = 0.1"

Answer. 0.1.